OpenStudy (anonymous):

Solve for the variable in the equation 3 (-2 - 3x) = -9x - 4 It is no solution right?

OpenStudy (anonymous):

\[-6-9x=-9x-4\] yes you are right

OpenStudy (angela210793):

-6-9x=-9x-4 -9x+9x=-4+6 0=2 no solution

OpenStudy (anonymous):

Thank you :D